OpenStudy (anonymous):

An evergreen nursery usually sells a certain shrub after 6 years of growth and shaping. The growth rate during those 6 years is approximated by dh/dt=3.5t+7 where t is the time in years and h is the height in centimeters. The seedlings are 15 centimeters tall when planted (t=0) . Find the height after t years. @phi

OpenStudy (stamp):

1. growth rate is approximated by dh/dt function. t = 6, so next we approximate the growth rate across the time span 2. find the growth rate dh/dt when t = 6. dh/dt = 28 cm / year 3. the plant is initially 15 cm, but add 28cm per year 4. evaluate function from step 3 at h(6). This is a bit of an odd question so a second opinion would be nice

OpenStudy (stamp):

@hayatabz1 make sure you look at the attached file do we have an answer key to this question, can we verify the result?

OpenStudy (anonymous):

thats not in the options... the options are @stamp

OpenStudy (anonymous):

a.11 b.23 c.47 d.52 e.29

OpenStudy (anonymous):

@stamp

OpenStudy (stamp):

idk man are you sure those answer choices are correct

OpenStudy (stamp):

all i can say at this point is find the height. cm = cm/year * years h = dh/dt * t

OpenStudy (stamp):

h = (3.5t + 7)*t h = 3.5t^2 + 7t

OpenStudy (stamp):

idfk

OpenStudy (anonymous):

oh my god sorry those were the wrong choicees these are the choices.. @stamp

OpenStudy (anonymous):

@stamp

OpenStudy (phi):

You have a differential equation \[ \frac{dh}{dt}= 3.5t+7 \] we can write this as \[ dh = 3.5 t + 7 dt\] now integrate both sides \[ \int dh = \int 3.5t + 7\ dt \] can you integrate this?

OpenStudy (anonymous):

yes 1.75t^2+6t+C?

OpenStudy (anonymous):

7t*

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

is C=15?

OpenStudy (anonymous):

so it would be h(t)=1.75^2+7t+15?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes 1.75t^2+6t+C? to be complete we get h + c1 = 1.75t^2+6t+c2 but c1 and c2 are unknown constants, so we can write this as h = 1.75t^2+6t+c2-c1 but c2-c1 is just an unknown constant, so just write it as h = 1.75t^2+6t+C (just like what you got, but with the gory details) to find C, replace t with 0 (because at t=0 we know h=15) 15= 0+0+C C= 15 now you have the full equation h= 1.75t^2+6t+15

OpenStudy (anonymous):

thank you very much you are writing 6t shouldnt it be 7t though? @phi

OpenStudy (phi):

yes, I meant 7t

OpenStudy (phi):

I copied and pasted your answer, and did not notice the 6t instead of 7 t

OpenStudy (phi):

*before you fixed it

OpenStudy (anonymous):

oh ok thank you @phi